Curriculum Vitae

Stephen G. Simpson

April 23, 2019

Stephen G. Simpson

Department of Mathematics 814-404-6176
Vanderbilt University  sgslogic@gmail.edu
Nashville, TN 37240, USA www.math.psu.edu/simpson/

Education:

1962–1966 Lehigh University, Bethlehem PA

B.A. (summa cum laude), M.S. in Mathematics

1966–1971 Massachusetts Institute of Technology, Cambridge MA

Ph.D. in Mathematics, thesis advisor Gerald E. Sacks

1969–1970 University of Wisconsin, Madison WI (no degree)

Scholarships, Fellowships, Research Grants, Awards:

1962–1966 Modern Transfer scholarship (Lehigh)
1966–1971 NSF and NDEA fellowships (M.I.T.)
1972–1974 NSF research grants (Yale, Berkeley)
1974–1975 Science Research Council research fellowship (Oxford)
1975–1996 NSF individual research grants (Penn State)
1980–1982 Alfred P. Sloan research fellowship (Penn State)
June 1981 CNRS research grant (Paris)
1983–1984 Deutsche Forschungsgemeinschaft research fellowship (Munich)
1986  Faculty Scholar Medal and prize (Penn State)
1987–1992 Raymond N. Shibley professorship (Penn State)
2000–2004 NSF individual research grant (Penn State)
2006–2010 NSF individual research grant (Penn State)
2007–2008 Grove Award for Interdisciplinary Research (Penn State)
2007–2011 NSF grant for collaborative research (Penn State)
2008–2009 Templeton Foundation research grant (Penn State)
2013–2019 Simons Foundation Collaboration Grant (Penn State, Vanderbilt)

Academic Employment:

1971–1972 Yale University (Gibbs Instructor)
1972–1974 University of California, Berkeley (Lecturer)
1974–1975 University of Oxford, England (Visiting Lecturer)
1975–1977 Pennsylvania State University (Assistant Professor)
1977–1980 Pennsylvania State University (Associate Professor)
Oct. 1978 University of Chicago (Visiting Associate Professor)
1979–1980 University of Connecticut (Visiting Associate Professor)

1980–2012  Pennsylvania State University (Professor)
June 1981  University of Paris, France (Visiting Professor)
1983–1984  University of Munich, Germany (Visiting Professor)
Jan.–Jun. 1987 Stanford University (Visiting Professor)
Jan.–Jun. 1992 University of Illinois (Visiting Professor)
1997–2000  University of Tennessee (Adjunct Professor)
Jan.–Jun. 1999 University of Tennessee (Visiting Professor)
2012–present Pennsylvania State University (Professor Emeritus)
2016–present Vanderbilt University (Research Professor)

Professional Activities:

Member of Phi Beta Kappa, 1965–present Member of Association for Symbolic Logic, 1970–present Member of American Mathematical Society, 1971–present Refereeing for various mathematical journals, 1971–present Reviewing for Mathematical Reviews, etc., 1972–present Co-editor of volume: Harvey Friedman’s Research in the Foundations of Mathematics, North-Holland Studies in Logic Series, 1985. Organizer of AMS Summer Research Conference “Applications of Mathematical Logic to Finite Combinatorics,” held August 4–10, 1985 at Humboldt State University, Arcata, California Member of Council, Association for Symbolic Logic, 1986–1989 Editor of volume: Logic and Combinatorics, Contemporary Mathematics, American Mathematical Society, 1987 Member of AMS/INS/SIAM Committee on Summer Research Conferences, 1987–1990 Member of Organizing Committee, 4th S.E. Asian Logic Conference, Tokyo, 1990 Chairman of Organizing Committee, spring meeting of the Association for Symbolic Logic (Penn State, April 7–8, 1990) Member of Committee on Conferences, Association for Symbolic Logic, 1990–1992. Associate Editor, Mathematical Logic Quarterly, 1992–2013. Program Committee, Fifth International Workshop on Termination, Utrecht, May 20–21, 2001. Editor of volume: Reverse Mathematics 2001, Association for Symbolic Logic, 2005, X + 401 pages. Member of Program Committee, Annual Meeting, Association for Symbolic Logic, Montreal, May 17–21, 2006. Member of Program Committee, Gödel Centennial Celebration, Association for Symbolic Logic, Montreal, May 2006.

Organizer, Special Session on Logic and Dynamical Systems, sponsored by the American Mathematical Society and the Association for Symbolic Logic as part of the Joint Mathematics Meetings, January 5–8, 2009, in Washington DC. Co-editor, Kurt Gödel: Essays for his Centennial, a volume of papers in honor of Gödel’s 100th birthday, published by the Association for Symbolic Logic and Cambridge University Press, 2010, VIII + 373 pages. Member of organizing committee, Computability Theory and Foundations of Mathematics, Tokyo, February 17–20, 2014. Co-chair of program committee, Computability Theory and Foundations of Mathematics, Tokyo, February 7–11, 2015. Member of program committee, New Challenges in Reverse Mathematics, Singapore, January 3–16, 2016. Attended 1-day conference for Simpson’s 70th birthday, University of Connecticut, May 22, 2016. Member of program committee, Computability Theory and Foundations of Mathematics, Singapore, September 8–12, 2017.

Invited Hour Addresses at International Meetings:

European Summer Meeting, Association for Symbolic Logic, Cambridge, England, 1971 Generalized Recursion Theory Symposium, Oslo, June 1972 Symposium in Mathematical Logic, Oslo, December 1974 European Summer Meeting, Association for Symbolic Logic, Clermont-Ferrand, France, July 1975 Spring Meeting, Association for Symbolic Logic, Chicago, April 1977 Symposium on Definability in Set Theory, Oberwolfach, West Germany, April 1977 Second Symposium on Generalized Recursion Theory, Oslo, June 1977 (short course of three lectures) Spring Meeting, Association for Symbolic Logic, Houston, April 1978 Kleene Symposium, Madison, Wisconsin, June 1978 European Summer Meeting, Association for Symbolic Logic, Leeds, England, August 1979 (short course of two lectures) Sixth International Congress for Logic, Methodology and Philosophy of Science, Hannover, West Germany, August 1979 Annual Meeting, Association for Symbolic Logic, New York, December 1979 European Summer Meeting, Association for Symbolic Logic, Prague, Czechoslovakia, August 1980 Logic Symposium, Patras, Greece, August 1980 Logic Meeting, M.I.T., April 1981 AMS Symposium on Recursion Theory, Cornell University, Ithaca, June–July 1972 Sixth Latin American Logic Symposium, Caracas, Venezuela (course of 10 lectures) Mathematical Logic Week, Oberwolfach, January 1984 Recursion Theory Week, Oberwolfach, April 1984 European Summer Meeting, Association for Symbolic Logic, Manchester, England, July 1984 Spring Meeting, Association for Symbolic Logic, Chicago, April 1985 Symposium on Hilbert’s Program Sixty Years Later, American Philosophical Association and Association for Symbolic Logic, Washington DC, December 1985 Symposium in Honor of Robin Gandy, Gregynog, Wales, September 1986 Workshop on Logic and Computer Science, Turin, Italy, October 1986 Logic Colloquium ’87, Granada, Spain, July 1987 Mathematical Society of Japan, Kyoto, October 1987 Third Southeast Asian Logic Conference, Beijing, China, October 1987 Interdisciplinary Conference on Axiomatic Systems, Columbus, Ohio, December 15–18, 1988 Logic Colloquium ’89, Berlin, July 1989 Conference on Set Theory and the Continuum Hypothesis, MSRI, Berkeley, October 1989. Fourth Southeast Asian Logic Conference, Tokyo, September 1990. Recursion Theory Conference, Oberwolfach, January 1996. Feferman Symposium, Stanford University, December 1998. Computability Theory and Applications, AMS Summer Conference, Boulder CO, June 1999. New York City Logic Conference, November 1999. Philosophy of Mathematics Workshop, CMU, December 8–9, 2000. Conference on Berechenbarkeitstheorie (“Computability Theory”), Oberwolfach, Germany, January 21–27, 2001. Special Session on Reverse Mathematics, ASL Annual Meeting, Philadelphia, March 10–13, 2001. Conference on Hilbert’s Problems Today, Pisa, Italy, April 5–7, 2001 Symposium on Reverse Mathematics and Computability Theory, ASL/APA Meeting, Minneapolis, May 3–5, 2001. Special Session on Proof Theory, AMS Meeting, Columbus, September 21–23, 2001. Annual Meeting, Association for Symbolic Logic, Las Vegas, June 1–4, 2002. Special Session on Computability and Models, American Mathematical Society, Baltimore, January 15–18, 2003.

Summer School and Workshop on Proof Theory, Computation and Complexity, 5 lectures, Technical University of Dresden, June 23–27, 2003. Mid-Atlantic Mathematical Logic Society, Long Island, NY, March 6, 2004. Special Session on Computability Theory and Applications, American Mathematical Society, Evanston, October 24, 2004. Special Session on Reverse Mathematics, AMS/ASL/MAA Joint Mathematics Meetings, Atlanta, January 7, 2005. Computational Prospects of Infinity, Institute of Mathematical Sciences, National University of Singapore, June 20 – August 15, 2005, two invited 1-hour talks. Workshop on Algorithmic Randomness, American Institute of Mathematics, Palo Alto, California, August 7–11, 2006. Special Session on Computability Theory (in honor of Manuel Lerman’s retirement), American Mathematical Society, Storrs, Connecticut, October 28–29, 2006. Topics in Computability (in honor of Richard Shore’s 60th birthday), Massachusetts Institute of Technology, January 21–22, 2007. Special Session on Computability Theory, Association for Symbolic Logic, Gainesville, Florida, March 10–13, 2007. Special Program in Proof Theory, Max Planck Institute for Mathematics, Bonn, Germany, March 1 – June 10, 2007. Workshop on Algorithmic Randomness, University of Chicago, September 15–19, 2007. Dynamical Systems Workshop, Pennsylvania State University, October 18–21, 2007. Workshop in Mathematical Logic, Mathematical Institute, Tohoku University, Sendai, Japan, February 20–23, 2008. Conference on Algorithmic Randomness, Nanjing University, Nanjing, China, May 19–23, 2008. Conference on Proof Theory and Constructive Mathematics, Mathematical Research Institute, Oberwolfach, Germany, April 6–12, 2008. Logic Colloquium 2008, Annual Summer Meeting, Association for Symbolic Logic, Bern, Switzerland, July 3–8, 2008. Conference on Reverse Mathematics, Banff International Research Station for Mathematical Innovation and Discovery, Banff, Canada, December 7–12, 2008. Special Session on Logic and Dynamical Systems, Washington DC, January 5–6, 2009. 15th Very Informal Gathering of Logicians (in honor of John Steel’s 60th birthday), UCLA, January 30 – February 1, 2009. Conference on Philosophy of Mathematics, New York University, April 3–5, 2009. Workshop on Algorithmic Randomness, University of Wisconsin, May 27–31, 2009.

11th Asian Logic Conference (in honor of C.-T. Chong), National University of Singapore, June 22–29, 2009. Reverse Mathematics: Foundations and Applications, University of Chicago, November 6–8, 2009. Workshop on Algorithmic Randomness, University of Hawaii, January 4–8, 2010. Conference on Dynamics and Computation, Centre International de Rencontres Mathématiques (CIRM), France, February 8–12, 2010. Fifth International Conference on Logic, Computability and Randomness, held at Notre Dame University, May 24–28, 2010. Workshop on Infinity and Truth, Institute of Mathematical Sciences, National University of Singapore, July 25–29, 2011. Reverse Mathematics Workshop, University of Chicago, September 16–18, 2011. Seventh International Conference on Computability, Complexity and Randomness, Isaac Newton Institute for Mathematical Sciences in Cambridge, England, July 2–6, 2012. Computability Theory and Foundations of Mathematics, Tokyo Institute of Technology, February 18–20, 2013. Sendai Logic School, Tohoku University, Japan, February 22–23, 2013. Computability and Complexity in Discrete and Continuous Worlds, American Mathematical Society at Iowa State University, April 27–28, 2013. Royal Society International Scientific Seminar, Computational Interpretations of Mathematical Theorems, Kavli Royal Society Centre, Milton Keynes, UK, November 25–26, 2013. Computability Theory and Foundations of Mathematics, Tokyo Institute of Technology, February 17–20, 2014. Computability, Complexity, and Randomness, Institute for Mathematical Sciences, National University of Singapore, June 9–13, 2014. Online Logic Seminar, Moscow State University, October 29, 2014 (via Skype). Conference on Proof Theory, Modal Logic and Reflection Principles, Instituto Tecnológico Autónomo de México (ITAM), Mexico City, September 29 – October 2, 2014. Scholar of Consequence, Annual Logic Lecture, University of Connecticut Group in Philosophical and Mathematical Logic, April 1–4, 2015. Workshop on Computability Theory, Bucharest, June 27–28, 2015. Computability in Europe, Bucharest, June 29 - July 3, 2015 (3-hour tutorial). Workshop on Reverse Mathematics, National University of Singapore, January 3–16, 2016. Public Lecture, National University of Singapore, January 6, 2016. Ordered Algebras and Logic, Vanderbilt University, February 6–7, 2016.

Algorithmic Randomness Interacts with Analysis and Ergodic Theory, Banff International Research Station, Casa Matemática Oaxaca, Mexico, December 4–9, 2016.

PhilMath Intersem 8, University of Paris 7, June 2–29, 2017.

Keynote address, Conference on Reverse Mathematics, University of Munich, October 9–13, 2017.

Conference on Axiomatic Thought, University of Lisbon, October 11–14, 2017.

Conference on Hermann Weyl’s monograph Das Kontinuum, University of Leeds, September 11–14, 2018.

Invited Talks at University Mathematics Departments:

Maryland, Berkeley, UCLA, Stanford, Texas, Penn State, Buffalo, Toronto, Manchester (England), Leeds (England), Cambridge (England), Oxford (England), Bedford College (U. of London, England), Cornell, Boston, Rockefeller, Connecticut, Chicago, M.I.T., Wisconsin, Paris (France), Lehigh, Bryn Mawr, McGill, Ohio State, Illinois at Champaign-Urbana, Yale, C.U.N.Y. Graduate Center, C.U.N.Y., Bowling Green, Duke, University of Florida, Florida State, I.V.I.C. (Caracas, Venezuela), Tübingen (Germany), Bielefeld (Germany), Munich (Germany), Utrecht (Netherlands), Münster (Germany), Hannover (Germany), Heidelberg (Germany), Berkeley, UCLA, Kobe (Japan), Nagoya (Japan), Kawai Institute (Japan), Tsukuba (Japan), Academia Sinica (Taiwan), Dickinson, Swarthmore, Torino (Italy), Notre Dame, Tennessee, National University of Singapore, University of Lisbon, Carnegie-Mellon, Notre Dame, University of Florida, University of Chicago, Maryland, Cornell, Moscow State (Russia, via Skype), Waseda (Japan), Vanderbilt, Louisiana State.

Invited Talks at Other University Departments:

Carnegie-Mellon (Computer Science, Philosophy), Columbia (Philosophy), University of Connecticut (Philosophy), Louisiana State (Philosophy). In addition Simpson has been active and given talks in the Mid-Atlantic Mathematical Logic Seminar (MAMLS), in the Logic Seminar, Computer Science Theory Seminar, and Mathematics Colloquium of the Pennsylvania State University, and in the Geometric Group Theory Seminar and the Universal Algebra and Logic Seminar at Vanderbilt University.

Publications:

More than 100 items. See separate publication list.

Ph.D. Theses Supervised:

1. John Steel, Determinateness and Subsystems of Analysis, Berkeley, 1977. (Steel is a tenured full professor at the University of California, Berkeley.)

[ Although Steel’s official thesis advisor was Professor John Addison of Berkeley, the following is a quotation from the acknowledgements page of Steel’s thesis. “I owe a great debt to Stephen Simpson, who guided me expertly in the perilous transition from study to research. The results of Chapters 1 and 2, together with less tangible aspects of my research, are a product of Simpson’s influence.” The thesis consists of three chapters. ]

2. Rick L. Smith, Theory of Profinite Groups with Effective Presentations, Pennsylvania State University, 1979. (Smith is a tenured associate professor at the University of Florida.)

3. Galen Weitkamp, Kleene Recursion over the Continuum, Pennsylvania State University, 1980. (Weitkamp is a tenured full professor at Western Illinois University.)

4. Peter Pappas, The Model Theoretic Structure of Group Rings, Pennsylvania State University, 1982. (Pappas is a tenured full professor and former department head at Vassar College.)

5. Stephen H. Brackin, On Ramsey-type Theorems and their Provability in Weak Formal Systems, Pennsylvania State University, 1984. (Brackin is a mathematician at Odyssey Research Associates.)

6. Mark Stephen Legrand, Coanalytic Sets in the Absence of Analytic Determinacy, Pennsylvania State University, 1985. (Legrand is an assistant professor at Auburn University.)

7. Douglas K. Brown, Functional Analysis in Weak Subsystems of Second Order Arithmetic, Pennsylvania State University, 1987. (Brown is a tenured professor at Livingstone College in North Carolina.)

8. Jeffry L. Hirst, Combinatorics in Subsystems of Second Order Arithmetic, Pennsylvania State University, 1987. (Hirst is a tenured professor at Appalachian State University in North Carolina.)

9. Xiaokang Yu, Measure Theory in Weak Subsystems of Second Order Arithmetic, Pennsylvania State University, 1987. (Connie Yu is a tenured professor at the New Jersey City University.)

10. Fernando Ferreira, Polynomial Time Computable Arithmetic and Conservative Extensions, Pennsylvania State University, 1988. (Ferreira is a tenured professor at the University of Lisbon.)

11. Kostas Hatzikiriakou, Commutative Algebra in Subsystems of Second Order Arithmetic, Pennsylvania State University, 1989. (Hatzikiriakou is a tenured instructor at the University of Thessalia, Greece.)

12. Alberto Marcone, Foundations of BQO Theory and Subsystems of Second Order Arithmetic, Pennsylvania State University, 1992. (Marcone is a tenured professor at the University of Udine, Italy.)

13. A. James Humphreys, On the Necessary Use of Strong Set Existence Axioms in Analysis and Functional Analysis, Pennsylvania State University, 1996. (Humphreys is a tenured instructor at Seattle University.)

14. Mariagnese Giusto, Topology, Analysis, and Reverse Mathematics, University of Torino, 1998. (Giusto is a postdoc at Notre Dame University.)

15. Stephen Binns, The Medvedev and Muchnik Lattices of Π_{1}^{0} Classes, Pennsylvania State University, 2003. (Binns is a professor at the University of Qatar.)

16. Carl Mummert, On the Reverse Mathematics of General Topology, Pennsylvania State University, 2005. (Mummert is a tenured professor at Marshall University.)

17. W. M. Phillip Hudelson, Scaled Randomness and Kolmogorov Complexity, Pennsylvania State University, 2013. (Hudelson is an associate at McKinsey Associates, Inc.)

18. Noopur Pathak, Computable Aspects of Measure Theory, Pennsylvania State University, 2013. (Pathak is employed by Ab Initio Software.)

19. Sankha S. Basu, A Model of Intuitionistic Higher-Order Logic Based on the Turing Degrees, Pennsylvania State University, 2013. (Basu is a professor at IIIT-Delhi, in India.)

20. Adrian Maler, Effective Theory of Lévy and Feller Processes, Pennsylvania State University, December 2015.

21. Hayden Jananthan, Vanderbilt University, 2021?